A company is attempting to determine the economic order quantity for glass used in the production of windows. 3,000 ponds of glass were used at a constant rate last period. Each order represents an ordering cost of $50. Carrying costs are $30 per pond over the 100-day planning period.

A-What is the economic order quantity?

B- The company has determined that it takes only 5 days to receive the order of glass after the placement of the order. When should the company order more glass?
Answer #1

A. EOQ =(2*100 day Demand*Ordering Cost/Carrying Cost)^0.5 =(2*3000*50/30)^0.5 =100

B, The level of reordering point =Average Daily Sales *Lead Time =(3000/100)*5 =150
When stock is 150 company should order more glass